OpenStudy (anonymous):

HELP PLEASE. Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -5 - 5 cos θ Origin only x-axis only y-axis only No symmetry

OpenStudy (anonymous):

@dan815 @phi @Hero @Miss.SweetiePie

OpenStudy (anonymous):

@paki

OpenStudy (campbell_st):

have you graphed the curve..?

OpenStudy (anonymous):

yes @campbell_st

OpenStudy (campbell_st):

is it polar coordinates..?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

I got x-axis only

OpenStudy (campbell_st):

well can you post your graph...?

OpenStudy (anonymous):

after replacing θ with -θ ... r = -5 - 5 cos(-θ) which means that both are even function so that cos(-θ) = cos(θ) r = -5 - 5 cos θ

OpenStudy (anonymous):

@Miss.SweetiePie thank u so much!!!

OpenStudy (campbell_st):

lol...

OpenStudy (anonymous):

Your welcome @em2000 are you a fan yet?

OpenStudy (michele_laino):

we can rewrite thhat equation, using the cartesian coordinates, namely: \[\Large r = - 5 - 5\frac{x}{r}\] from which we get: \[\Large {x^2} + {y^2} + 5\sqrt {{x^2} + {y^2}} + 5x = 0\] Now it is easy to understand if there is a symmetry for that equation